OpenStudy (anonymous):

What polynomial has a graph that passes through the given points? (-2, 2) (-1, -1) (1, 5) (3, 67)

OpenStudy (amistre64):

simple enough

OpenStudy (amistre64):

\[y = c_0+c_1(x-x_1)+c_2(x-x_1)(x-x_2)+c_2(x-x_1)(x-x_2)(x-x_3)\]

OpenStudy (amistre64):

\[y = c_0+c_1(x+2)+c_2(x+2)(x+1)+c_2(x+2)(x+1)(x-1)\] x=-2; c0=2 \[y = 2 +c_1(x+2)+c_2(x+2)(x+1)+c_2(x+2)(x+1)(x-1)\]each constant term gets filled in by progressing to the next point that zeros out the unknowns

OpenStudy (amistre64):

x=-1, y=-1 \[-1 = 2+c_1(-1+2)+0\] \[1 = c_1(1)\] c1=1

OpenStudy (amistre64):

determine the other constants in this manner, then expand it all out as desired

OpenStudy (amistre64):

c1=-3 tho :) forgot how to add

OpenStudy (dumbcow):

you can use excel to check your work

OpenStudy (amistre64):

someones getting a bonus check from bill gates hmmmm :)

OpenStudy (dumbcow):

haha :)